OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9

OpenStudy (anonymous):

From First equation : \[y = 4x - 8\] Put it in second equation, we have \[2x + 3(4x - 8) = 9\] \[2x + 12x - 24 = 9\] \[14x = 9+24\] \[x = 33/14\] And therefore, y = 10/7.. Understood or not???